OpenStudy (anonymous):

I have a question regarding absolute value: |x| = \sqrt{(x)^2} If this is the case, then why doesn't the square and the square root just cancel each other out, leaving you with: |x| = x ?

OpenStudy (jamesj):

Put in x = -1 and you will see why: \[x^{2} = (-1)^2 = 1 \ \hbox{hence} \ \sqrt{x^2} = \sqrt{1} = 1.\]

OpenStudy (anonymous):

I don't understand.

OpenStudy (jamesj):

Your argument is that |x| = x for every real number x. Using the example of x = -1 you can see that this is in fact not the case, as \[|x| = |-1| = 1\] but x = -1. Hence it is not the case for x = -1 that |x| = x. And therefore it is not the case in general that |x| = x for all x.

myininaya (myininaya):

\[\sqrt{x^2}=x if x>0\] \[\sqrt{x^2}=-x if x<0\] \[\sqrt{x^2}=0 if x=0\]